OpenStudy (anonymous):

what is the simplified form of: 7 radical X^5 times 7 radical X^5?

OpenStudy (anonymous):

the choices are (A) radical x (B) x 4/9 (C) x (D) x^7 radical x^3

ganeshie8 (ganeshie8):

\(\large \sqrt[7]{X^5} \times \sqrt[7]{X^5}\)

ganeshie8 (ganeshie8):

is the question like this ?

OpenStudy (anonymous):

yes. @ganeshie8

ganeshie8 (ganeshie8):

use this, a x a = a^2

ganeshie8 (ganeshie8):

\(\large \sqrt[7]{X^5} \times \sqrt[7]{X^5} = (\sqrt[7]{X^5} )^2\)

OpenStudy (anonymous):

thank you

ganeshie8 (ganeshie8):

\(\large \sqrt[7]{X^{10}} \)

ganeshie8 (ganeshie8):

\(\large \sqrt[7]{X^{7} \times X^3} \)

ganeshie8 (ganeshie8):

\(\large X \sqrt[7]{X^3} \)

ganeshie8 (ganeshie8):

see if that makes some sense :)

OpenStudy (anonymous):

it does :)

ganeshie8 (ganeshie8):

glad it does :D